OpenStudy (anonymous):

2 3+a+ 2 -- - --- + --------- = a+3 2a+6 a²+6a+9a² solve the equation

OpenStudy (anonymous):

notice that 2a+6 = 2(a+3) and a^2+6a+9 = (a+3)^2

OpenStudy (anonymous):

2 3+a 2 -- - --- + --------- = a+3 2a+6 a²+6a+9a² Solve